OpenStudy (anonymous):

Hi, anyone please recommend best explanatory material for the D operator method of solving ordinary differential equations?

jimthompson5910 (jim_thompson5910):

D(some function) = that function's derivative Ex: f(x) = x^2+9x D( f(x) ) = D( x^2+9x ) = 2x+9

jimthompson5910 (jim_thompson5910):

so instead of saying \[\Large \frac{dy}{dx}\], you can shorten it and just say D (keep the context in mind though)

OpenStudy (zarkon):

one can also add a subscript to indicate the variable you are differentiating with respect to... \[D_x(f(x))=\frac{d}{dx}f(x)=f'(x)\]